z/z +3 = 3z/5z - 1
Answer by HyperBrain(694)

z/z +3 = 3z/5z - 1


By cross multiplication,
z(5z-1)=3z(z+3)
5z2-z=3z2+9z
2z2-10z=0
z2-5z=0
z(z-5)=0
z=0 or z=5
Power up,
HyperBrain!
Bored? Go to my blog.
hyperclean